# A Framework for Approximating Perturbed Optimal Control Problems

### v1

@@ 1 Introduction
Section 2 outlines our methods in generality, allowing for its application to a wide range of problems. Our first goal is to reduce the number of significant parameters using global sensitivity analysis. Then, rather than totally recomputing a solution, we focus on approximating a solution, generalizing a method presented in [1]. By storing and interpolating sensitivity information, which requires less memory due to parameter reduction, we may further reduce in-transit computation time without significant sacrifices of accuracy. These three steps, (i) parameter reduction, (ii) sensitivity information interpolation, and (iii) approximating using a general method from [1], individually constitute ways to improve computation time given uncertainty, and we consider the efficacy of each.
Section 3 applies this solution method to a model Space Shuttle Trajectory problem from [4]. There we solve for the optimal control angle which maximizes the longitude travelled in a model space shuttle’s reentry. This provides an in-depth example that demonstrates the strengths and limitations of each part of the procedure.
@@ 2.2 Dimension Reduction
for the parameters. We wish to compute Sobol indices, which measure the amount of output variance that is caused by variance in input parameters. While Sobol indices provide a clear ranking of how important parameters are, computing them can be expensive. Therefore, we bound Sobol indices using derivative-based global sensitivity measures (DGSMs) as described in [6]. Let $\boldsymbol{u}:\Theta\rightarrow\mathbb{R}^{N+1}$ be a vector-valued function. The derivative-based global sensitivity measure of $\boldsymbol{u}$ with respect to parameter $j$ is defined as
@@
Numerically, the DGSMs may be computed via a Monte Carlo method and the terms $\frac{\partial u_{i}}{\partial\theta_{j}}$ may computed via Hyper-Differential Sensitivity Analysis as seen in Section 2.2.1.
In particular, a quasi-Monte Carlo (QMC) method employing low-discrepancy sequences to sample points, as presented in [9], is shown to converge faster than a standard Monte Carlo method when computing DGSMs. It is also possible to vectorize the process by interchanging the QMC summation and DGSM summation. With this interchange we have,
@@ 2.3.1 Generalizing an Approximation Method
so $\boldsymbol{u}^{*}(\boldsymbol{\theta_{1}})\approx\boldsymbol{u}^{*}_{M}$. Since this method does require re-computation of the derivative at various points, it trades accuracy for some computational speed. This is well-known to have a local truncation error proportional to the square of the step size, or $\mathcal{O}(||\boldsymbol{\theta_{1}}-\boldsymbol{\theta_{0}}||_{2}^{2})$. In the case of in transit changes or large perturbations, this accuracy becomes extremely important. However, due to compounding error in calculating derivatives and steps, the number of step sizes that improve the approximation is limited by the system involved.
Any number of other numerical differential equation methods may be applied instead of a forward-Euler method, such as higher order Runge-Kutta methods which may help control when the ODE acts stiff at the cost of computation time.
@@ 2.3.2 Interpolation of the Key Jacobian
Ironically, an increased number of steps tends to outweigh higher order methods in terms of accuracy in practice. While several other ODE methods, such as Runge Kutta 3, Adams-Bashforth multistep, Adams-Moulton multistep, and Gaussian Quadrature, were considered, we will focus on the forward Euler method moving forward and show the effectiveness of the interpolated step in section 3.
While it is not implemented in this paper, storing higher-order derivatives to approximate the Jacobian $\frac{\partial u_{i}}{\partial\theta_{j}}$ could similarly improve computational efficiency without sacrificing much accuracy.
@@ 3 Numerical Results
3:  Compute DGSMs (7) and reduce the parameter space using the Sobol Indice upper bounds (8)
@@ 3.1.1 Problem Formulation
(18) $\boldsymbol{x}(0)=\begin{cases}h(0)&=80000\\ v(0)&=2500\\ \gamma(0)&=-5\frac{\pi}{180}\end{cases}$
@@ 3.1.2 Parameter Space Reduction
Parameter # DGSM Upper Bound
$m$ 0.00322 0.0213
$\rho_{0}$ 0.00356 0.0235
$a_{0}$ 0.00991 0.0654
$a_{1}$ 2.66 17.6
$b_{0}$ 0.656 4.33
$b_{1}$ 0.00539 0.0356
$b_{2}$ 0.750 4.95
@@ 3.1.3 Space Shuttle Results
Figure 1: Controller, $\boldsymbol{u}$, and state variables, $\boldsymbol{x}$, of the optimal solution and approximated solution for parameter shift $\boldsymbol{\theta}_{0}\to\boldsymbol{\theta}_{1}$ at $t=2000$.
@@
In Figure 2 we see that both the cost and the norm error skew right, thus implying most interpolated step approximations efficiently minimize cost and accurately approximate the optimal solution for the parameter change. We chose these two error metrics because we noticed that an approximation that minimizes error does not necessarily minimize cost and vice versa. Physically, this implies there are multiple paths that lead to similar results, and computationally this implies that this is a highly non-convex problem with many local minima. By using both metrics they give a more complete view of the effectiveness of our approximation.
@@
Mean 1.3729 1.1104 0.0621
Median -1.6338 -1.5973 0.0211
where $||u||$ is on the scale of $0.68$. The median difference represents about a $3\%$ change in the controller norm and an even smaller change in the optimal solution’s cost. Thus, optimizing with the reduced parameters returns solutions with nearly the exact same cost and almost no difference in path.
@@
Mean 1.3729 $2.54\times 10^{3}$ 683.1728
Median -1.6338 -0.9683 -0.9419
@@
$\left\lVert u_{lin}^{7}-u_{lin}^{3}\right\rVert$ $\left\lVert u_{opt}^{7}-u_{lin}^{7}\right\rVert$ $\left\lVert u_{opt}^{7}-u_{lin}^{3}\right\rVert$ $\left\lVert u_{opt}^{3}-u_{lin}^{7}\right\rVert$ $\left\lVert u_{opt}^{3}-u_{lin}^{3p}\right\rVert$
Mean 0.0427 0.1480 0.1447 0.1443 0.1371
Median 0.0437 0.1212 0.1187 0.1260 0.1165
Table 3 provides two main results. First, the cost difference between an approximation using all parameters versus one using the reduced parameters is minuscule, around $3\%$. This suggests that parameter reduction is an effective method for reducing computation cost while maintaining accuracy. Secondly, we see that on average a linear approximation does not do a good job of minimizing cost and has many outliers that greatly reduce its accuracy.
Table 4 provides similar results. We see that the norm difference between approximations using the full parameter space versus the reduced space is extremely small. This once again suggests that the parameter space reduction does not reduce accuracy of approximations. Secondly, we see that error between any optimal solution and linear approximation is large, once again showing that a linear approximation is not effective at approximating an optimal solution.
@@
Mean 1.3729 1.8242 1.3374
Median -1.6338 -1.4618 -1.3891
@@
Mean 0.0756 0.1103 0.1152 0.1196 0.1038
Median 0.0463 0.0590 0.0542 0.0700 0.0470
Table 5 shows a similar narrative to the case without interpolation. First, we see that the mean and median are distant because as mentioned earlier, imposing constraints as penalties represents itself as discrepancy. However, this has greatly improved the stability of the result compared to the linear case. The mean and median costs now closely resemble that of the optimized results showing that our approximation method does a great job of minimizing cost, just as seen and discussed for figure 2. Finally, we once again see that the minimal difference between the 7- and 3-parameter cases justifies the use of parameter reduction as it also greatly reduces the computational load of using the interpolation.
Table 6 also shows similar norm error results to the linear case. The norm difference between the 7- and 3-parameter approximations is small enough to again imply that parameter reduction is an effective method for reducing computation cost while maintaining accuracy. Additionally, we see much lower errors showing that the interpolated step approximation method accurately replicates the optimal solution while only using precomputed data.
@@ 4 Conclusion
The space shuttle example of Betts verifies the parameter reduction and interpolated-step approximation methods. While this is a limited application, the generality of these methods allows their accuracy to be controlled by the user and allows for more complex application. Stricter methods of applying constraints would also likely limit the instability possible in the approximated solution. The proposed approach has applications in aerospace trajectory control, parameter uncertainty calibration, and broad applicability to optimal control problems in which parameter space reduction could improve computation time with little to no reduction in accuracy.
@@ Appendix B Space Shuttle Cost Function
Given below is the complete cost function for the 2 DOF space shuttle problem. The first term aims to maximize longitudinal distance. The next three penalties enforce terminal conditions and the final four penalties enforce state constraints. The $\beta_{i}$ are tuning parameters.

### v2

@@ 1 Introduction
Section 2 outlines our methods in generality, allowing for its application to a wide range of problems. Our first goal is to reduce the number of significant parameters using global sensitivity analysis. Then, rather than totally recomputing a solution, we focus on approximating a solution, generalizing a method presented in [1]. By storing and interpolating sensitivity information, which requires less memory due to parameter reduction, we may further reduce in-transit computation time without significant sacrifices of accuracy. These three steps, (i) parameter reduction, (ii) sensitivity information interpolation, and (iii) approximating using a general method from [1], individually constitute ways to improve computation time given uncertainty, and we consider the efficacy of each. Hence, our approach is to adapt the optimal trajectory dynamically based on sensitivity information. Typical feedback control which relies on reactive adjustments, works well for small deviations but may struggle with large deviations, and so this new approach can handle those larger perturbations given its adjustments made using precomputed sensitivity information.
Thus, we want to emphasize that our method is intended to complement, rather than replace, existing feedback control techniques, which remain well-suited for handling small deviations. Additionally, this approach represents a singular strategy within optimal control, and its applicability should be evaluated based on the specific problem at hand. Users should consider its usefulness in conjunction with other established methods to determine the most effective solution for their particular application.
Section 3 applies this solution method to a model Space Shuttle Trajectory problem from [4]. There we solve for the optimal control angle which maximizes the longitude traveled in a model space shuttle’s reentry. This provides an in-depth example that demonstrates the strengths and limitations of each part of the procedure.
@@ 2.2 Dimension Reduction
for the parameters. We wish to compute Sobol indices, which measure the amount of output variance that is caused by variance in input parameters [10]. It is important to note that Sobol indices come in three different forms, main, interaction, and total. The main Sobol index is simply the ratio of the variance of a single input, say the parameter $\theta_{j}$, by the total variance of the output, in our case the optimal controller $\boldsymbol{u}$. Thus, this main index measures the direct effects of the input variable and does not account for any interaction effects with other variables. These interaction indices capture how the combined variations of input parameters like $\theta_{i}$ and $\theta_{j}$ affect the model output. The total Sobol index captures the contribution of an input variable to the output variance, accounting for both its main effects and all interaction effects involving the input variable.
These total Sobol indices provide a clear ranking of how important input parameters are. Total Sobol indices range from 0 to 1, yet will not sum to 1 given the overlapping of interaction terms. These indices would allow the most unimportant parameters to be reduced to their nominal values and ignored for perturbation analysis. However, computing them can be expensive. Therefore, we bound Sobol indices using derivative-based global sensitivity measures (DGSMs) as described in [6].
Note that, if Sobol indices could be precisely computed, the main Sobol indices would provide a more valid reduction of parameter space. This is because, when setting a reduced parameter to its nominal value, we are then ignoring all interaction effects, leaving only the direct effects captured by the main Sobol indices. As such, the main Sobol indices would better describe how important each parameter is and whether it should be retained in the model.
However, since we are working with upper bounds on Sobol indices, the total index also provides a valid measure of whether or not we should retain or discard a parameter. The total index accounts for both direct and interaction effects, providing a comprehensive measure of a parameter’s maximum possible contribution to output variance. For example, given a input parameter with a small upper bound on its total index, we can then safely say that even its maximum possible contribution (direct and interaction effects) to the output variance is negligible.
Nevertheless, the use of upper bounds comes with limitations. While we can confidently identify parameters with negligible total contributions, if given a large upper bound, i.e. an upper bound above 1, we cannot infer the relative importance of the parameter and its direct versus interaction effects. Thus, we leave these parameters with unclear significance to be further analyzed through the perturbation-based methods in our framework.
The structure of computing the upper bounds on total Sobol indices is as follows. Let $\boldsymbol{u}:\Theta\rightarrow\mathbb{R}^{N+1}$ be a vector-valued function. The derivative-based global sensitivity measure of $\boldsymbol{u}$ with respect to parameter $j$ is defined as
@@
Numerically, the DGSMs may be computed via a Monte Carlo method and the terms $\frac{\partial u_{i}}{\partial\theta_{j}}$ may computed via Hyper-Differential Sensitivity Analysis as seen in Section 2.2.1. In particular, a quasi-Monte Carlo (QMC) method employing low-discrepancy sequences to sample points, as presented in [9], is shown to converge faster than a standard Monte Carlo method when computing DGSMs. It is also possible to vectorize the process by interchanging the QMC summation and DGSM summation. With this interchange we have,
@@ 2.3.1 Generalizing an Approximation Method
so $\boldsymbol{u}^{*}(\boldsymbol{\theta_{1}})\approx\boldsymbol{u}^{*}_{M}$. Since this method does require re-computation of the derivative at various points, it trades accuracy for some computational speed. This is well-known to have a local truncation error proportional to the square of the step size, or $\mathcal{O}(||\boldsymbol{\theta_{1}}-\boldsymbol{\theta_{0}}||_{2}^{2})$. In the case of in transit changes or large perturbations, this accuracy becomes extremely important. However, due to compounding error in calculating derivatives and steps, the number of step sizes that improve the approximation is limited by the system involved. Any number of other numerical differential equation methods may be applied instead of a forward-Euler method, such as higher order Runge-Kutta methods which may help control when the ODE acts stiff at the cost of computation time.
@@ 2.3.2 Interpolation of the Key Jacobian
Ironically, an increased number of steps tends to outweigh higher order methods in terms of accuracy in practice. While several other ODE methods, such as Runge Kutta 3, Adams-Bashforth multistep, Adams-Moulton multistep, and Gaussian Quadrature, were considered, we will focus on the forward Euler method moving forward and show the effectiveness of the interpolated step in section 3. Additionally, while it is not implemented in this paper, storing higher-order derivatives to approximate the Jacobian $\frac{\partial u_{i}}{\partial\theta_{j}}$ could similarly improve computational efficiency without sacrificing much accuracy.
@@ 3 Numerical Results
3:  Compute DGSMs (7) and reduce the parameter space using the Sobol Index upper bounds (8)
@@ 3.1.1 Problem Formulation
(18) $\boldsymbol{x}(0)=\begin{cases}h(0)&=260000\\ v(0)&=25600\\ \gamma(0)&=-\frac{\pi}{180}\end{cases}$
@@ 3.1.2 Parameter Space Reduction
{tblr}
colspec=——c c c——, rowsep=3pt, hlines, row6=gray!30, row5=gray!30, row8=gray!30 Parameter # & DGSM Upper Bound
$m$ $3.22\times 10^{-3}$ $2.13\times 10^{-2}$
$\rho_{0}$ $3.56\times 10^{-3}$ $2.35\times 10^{-2}$
$a_{0}$ $9.91\times 10^{-3}$ $6.54\times 10^{-2}$
$a_{1}$ $2.66\times 10^{0}$ $1.76\times 10^{1}$
$b_{0}$ $6.56\times 10^{-1}$ $4.33\times 10^{0}$
$b_{1}$ $5.39\times 10^{-3}$ $3.56\times 10^{-2}$
$b_{2}$ $7.50\times 10^{-1}$ $4.95\times 10^{0}$
@@ 3.1.3 Space Shuttle Results
Figure 1: Controller, $\boldsymbol{u}$, and state variables, $\boldsymbol{x}$, of the optimal solution and approximated solution for parameter shift $\boldsymbol{\theta}_{0}\to\boldsymbol{\theta}_{1}$ at $t=2000$. The terminal constraints are represented by the thin dashed lines.
@@
In Figure 2 we see that both the cost and the norm error skew right, thus implying most interpolated step approximations efficiently minimize cost and accurately approximate the optimal solution for the parameter change. We chose these two error metrics because we noticed that an approximation that minimizes error does not necessarily minimize cost and vice versa. Since we only have control over the costs here, this implies there are multiple paths that lead to similar results, and computationally this implies that this is a highly non-convex problem with many local minima. By using both metrics, a more complete view of the effectiveness of our approximation is shown.
@@
Finally, Figure 2 displays a bimodal distribution with a main mode around 0 and a second mode around 0.2. We notice that this secondary mode seems to be a feature of the problem. As more samples are taken, the tail extends and the mode persists. However, we can calculate,
$\mathbb{P}\left(\lVert u_{opt}-u_{IS}\rVert<0.2\right)$ $=0.84.$
Thus, there is an 84% chance that our approximation is below this secondary mode and thus is an excellent error approximation. Additionally, given that we are looking at a considerable 10% perturbation range, we would expect this probability to only increase with a more local approximation perturbation range of 5% or less.
@@
Mean $1.37\times 10^{0}$ $1.11\times 10^{0}$ $6.21\times 10^{-2}$
Median $-1.63\times 10^{0}$ $-1.60\times 10^{0}$ $2.11\times 10^{-2}$
where $||u||$ is on the scale of $0.68$. The median difference, 0.0211, represents about a $3\%$ change in the controller norm and an even smaller change in the optimal solution’s cost. Thus, optimizing with the reduced parameters returns solutions with nearly the exact same cost and almost no difference in path.
@@
Mean $1.37\times 10^{0}$ $2.54\times 10^{3}$ $6.83\times 10^{2}$
Median $-1.63\times 10^{0}$ $-9.68\times 10^{-1}$ $-9.42\times 10^{-1}$
@@
$\left\lVert u_{lin}^{7}-u_{lin}^{3}\right\rVert$ $\left\lVert u_{opt}^{7}-u_{lin}^{7}\right\rVert$ $\left\lVert u_{opt}^{7}-u_{lin}^{3}\right\rVert$ $\left\lVert u_{opt}^{3}-u_{lin}^{7}\right\rVert$ $\left\lVert u_{opt}^{3}-u_{lin}^{3}\right\rVert$
Mean $4.27\times 10^{-2}$ $1.48\times 10^{-1}$ $1.45\times 10^{-1}$ $1.44\times 10^{-1}$ $1.37\times 10^{-1}$
Median $4.37\times 10^{-2}$ $1.21\times 10^{-1}$ $1.19\times 10^{-1}$ $1.26\times 10^{-1}$ $1.17\times 10^{-1}$
Table 3 provides two main results. First, the difference in median values of $J(u_{\text{lin}}^{7})$ and $J(u_{\text{lin}}^{3})$, i.e. $-9.68\times 10^{-1}$ and $-9.42\times 10^{-1}$ respectively, is minuscule, around a $3\%$ difference. This suggests that parameter reduction is an effective method for reducing computation cost while maintaining accuracy. Secondly, we see that on average a linear approximation does not do a good job of minimizing cost and has many outliers that greatly reduce its accuracy. This is shown by the high mean costs of $2.54\times 10^{3}$ and $6.83\times 10^{2}$.
Table 4 provides similar results. We see that the norm difference between approximations using the full parameter space versus the reduced space is extremely small, $4.37\times 10^{-2}$. This once again suggests that the parameter space reduction does not reduce accuracy of approximations. However, we see that error between any optimal solution and linear approximation is large, such as $\left\lVert u_{opt}^{7}-u_{lin}^{7}\right\rVert$ median of $1.21\times 10^{-1}$ which is an 82% increase from the nominal normed controller. This once again shows that a linear approximation is not effective at approximating an optimal solution.
@@
Mean $1.37\times 10^{0}$ $1.82\times 10^{0}$ $1.34\times 10^{0}$
Median $-1.63\times 10^{0}$ $-1.46\times 10^{0}$ $-1.39\times 10^{0}$
@@
Mean $7.56\times 10^{-2}$ $1.10\times 10^{-1}$ $1.15\times 10^{-1}$ $1.20\times 10^{-1}$ $1.04\times 10^{-1}$
Median $4.63\times 10^{-2}$ $5.90\times 10^{-2}$ $5.42\times 10^{-2}$ $7.00\times 10^{-2}$ $4.70\times 10^{-2}$
Table 5 shows a similar narrative to the case without interpolation. First, we see that the mean and median are distant because as mentioned earlier, imposing constraints as penalties represents itself as discrepancy. However, this has greatly improved the stability of the result compared to the linear case. The mean and median costs now closely resemble that of the optimized result, showing that our approximation method does a great job of minimizing cost, just as seen and discussed for figure 2. Finally, we once again see that the minimal difference between the 7- and 3-parameter cases, median costs of $-1.46$ and $-1.39$ respectively, justifies the use of parameter reduction as it also greatly reduces the computational load of using the interpolation.
Table 6 also shows similar norm error results to the linear case. The norm difference between the 7- and 3-parameter approximations, $\left\lVert u_{IS}^{7}-u_{IS}^{3}\right\rVert$ having a median of $4.63\times 10^{-2}$, is small enough to again imply that parameter reduction is an effective method for reducing computation cost while maintaining accuracy. Additionally, we see much lower errors showing that the interpolated step approximation method accurately replicates the optimal solution while only using precomputed data.
@@ 3.2 Space Shuttle Discussion
We observe a significant reduction in computation time, as shown in Table 7. As mentioned earlier, recalculating the optimal solution in-transit is often prohibitive given the lack of any guarantee of a favorable solution in any short period of time.
Table 7: 500 Perturbed Parameters Approximations Mean Computation Times
$\text{time}(u_{opt})$ $\text{time}(u_{IS})$
Mean (s) $3.19\times 10^{1}$ $1.12\times 10^{-1}$
Table 7 highlights the considerable decrease in computation time between obtaining the optimal solution and our interpolated step approximation. Note that the primary factor that is limiting the approximation from achieving even faster computation times is the speed of the interpolation function in matlab. Additionally, keep in mind that the data in Table 7 was acquired on a local workstation as opposed to the specialized chips and hardware that would be onboard this aircraft problem.
@@ 4 Conclusion
As stated in the introduction, we want to emphasize that our method is intended to be used in conjunction with established methods for handling small parameter perturbations such as feedback control. The space shuttle example of Betts verifies the parameter reduction and interpolated-step approximation methods. While this is a limited application, the generality of these methods allows their accuracy to be controlled by the user and allows for more complex application. Stricter methods of applying constraints would also likely limit the instability possible in the approximated solution. It is important to note that this approach is one of many in the toolbox for solving perturbation issues in optimal control problems and thus users should consider other established methods alongside this new approach to determine the most effective and practical solution for the application under consideration. Yet, the proposed approach has applications in aerospace trajectory control, parameter uncertainty calibration, and broad applicability to optimal control problems in which parameter space reduction could improve computation time with little to no reduction in accuracy.
@@ Appendix B Space Shuttle Cost Function
Given below is the complete cost function for the 2 DOF space shuttle problem. The first term aims to maximize longitudinal distance. The next three penalties enforce terminal conditions and the final four penalties enforce state constraints. The $\beta_{i}$ are tuning parameters. These tuning parameter values were chosen experimentally and are going to be application dependent.